OpenStudy (anonymous):

negative 8 to the five-thirds power can be simplified to -32.

OpenStudy (anonymous):

True or false

OpenStudy (anonymous):

True

OpenStudy (yttrium):

Let's check. (-8)^(5/3) = cuberoot(-8)^5 = -2^5 = -32 Therefore?